suppose the prices of all three year old porsche 911 sports cars have a normal distribution with a mean of 48,125 and a standard deviation of 1600. what is the probability that a randomly selected three year old porsche 911 will sell for a price between 46,000 and 49,000?
Answer by stanbon(75887)

suppose the prices of all three year old porsche 911 sport)s cars have a normal distribution with a mean of 48,125 and a standard deviation of 1600. what is the probability that a randomly selected three year old porsche 911 will sell for a price between 46,000 and 49,000?
-----------------
Determine the z-value of 46000 and of 49000
z(46000) = (46000-48125)/1600 = -1.328125..
z(49000) = (49000-48125)/1600 = 875/1600 = 0.546875...
-------------
P(46000 < x < 49000) = P(-1.328125 < z < 0.546875..) = 0.6157
================
Cheers,
Stan H.